OpenStudy (highschoolmom2010):

A triangle has side lengths 10, 15, and 7. Is the triangle acute, obtuse, or right? Explain.

OpenStudy (anonymous):

acute angled triangle

OpenStudy (anonymous):

If A^2 + B^2 = C^2 is satisfied, it is a right triangle. If A^2 + B^2 > C^2 is satisfied, it an acute triangle. If A^2 + B^2 < C^2 is satisfied, it an obtuse triangle. Where A and B are any 2 sides that isn't the longest and C is the longest side. For #1, 5^2 + 2^2 _ 6^2 25 + 9 _ 36 34 < 36 So it is an obtuse triangle. You can use this for the other two as well.

OpenStudy (highschoolmom2010):

i put acute but my teacher said that it was incorrect...

OpenStudy (highschoolmom2010):

\[10^2+15^2=7^2\] \[100+225\neq49\] \[325>49\]

OpenStudy (highschoolmom2010):

unless i am doing it wrong.....

ganeshie8 (ganeshie8):

u made a mistake in choosing sides

ganeshie8 (ganeshie8):